Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1
Answer by nerdybill(7384)

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1
.
vertex is mid-way between the focus and directrix:
vertex: (-5,2)
standard form of a vertical parabola
(x – h)^2 = 4p(y – k)
(x – (-5))^2 = 4p(y – 2)
(x + 5)^2 = 4p(y – 2)
.
p is the distance between the vertex and either the focus or directrix:
p = 3
.
so our final equation is:
(x + 5)^2 = 4*3(y – 2)
(x + 5)^2 = 12(y – 2)